# Simple Domain Adaptation for Sparse Retrievers

### v1

@@ Abstract
In Information Retrieval, and more generally in Natural Language Processing, adapting models to specific domains is conducted through fine-tuning. Despite the successes achieved by this method and its versatility, the need for human-curated and labeled data makes it impractical to transfer to new tasks, domains, and/or languages when training data doesn’t exist. Using the model without training (zero-shot) is another option that however suffers an effectiveness cost, especially in the case of first-stage retrievers. Numerous research directions have emerged to tackle these issues, most of them in the context of adapting to a task or a language. However, the literature is scarcer for domain (or topic) adaptation. In this paper, we address this issue of cross-topic discrepancy for a sparse first-stage retriever by transposing a method initially designed for language adaptation. By leveraging pre-training on the target data to learn domain-specific knowledge, this technique alleviates the need for annotated data and expands the scope of domain adaptation. Despite their relatively good generalization ability, we show that even sparse retrievers can benefit from our simple domain adaptation method. [^1] This paper is a preprint that was accepted at ECIR 2024
@@ 4.1 Datasets, baselines, and ablations
Dataset ($\downarrow$) BM25 0-shot
0 layer
1 layer
2 layers
4 layers
w/o pre-training w/o source
TREC-COVID 58.1 71.3
$68.8^{\dagger}$
$70.5^{\dagger}$
$\textbf{72.1}^{\dagger}$
$70.9^{\dagger}$
$67.9^{\dagger}$ $67.7^{\dagger}$
NFCorpus 24.5 24.9 25.6 25.4 25.1
24.1
25.1 25.6
BioASQ 52.3 43.1
$\textbf{45.4}^{\ddagger}$
44.2
$\textbf{45.2}^{\ddagger}$
$\textbf{46.2}^{\ddagger}$
$\textbf{44.8}^{\ddagger}$ $\textbf{45.3}^{\ddagger}$
FiQA-2018 23.6 27.3
$27.2^{\dagger}$
$\textbf{27.5}^{\dagger}$
$\textbf{29.5}^{\dagger\ddagger}$
$\textbf{29.0}^{\dagger\ddagger}$
$25.2^{\dagger}$ $\textbf{28.3}^{\dagger}$
TREC-NEWS 31.4 32.1 33.3 32.8 33.0
$\textbf{36.5}^{\ddagger}$
31.9 33.1
Robust04 40.8 39.6
$\textbf{42.9}^{\ddagger}$
$\textbf{42.9}^{\ddagger}$
$\textbf{43.5}^{\ddagger}$
$\textbf{42.6}^{\ddagger}$
41.2 $\textbf{42.1}^{\ddagger}$
ANTIQUE 45.4 43.3
42.9
44.0 43.7 44.1 $\textbf{45.0}^{\ddagger}$ $\textbf{45.0}^{\ddagger}$
LoTTE-Rec. 39.3 46.2
$\textbf{47.6}^{\dagger\ddagger}$
$\textbf{47.7}^{\dagger\ddagger}$
$\textbf{47.8}^{\dagger\ddagger}$
$\textbf{46.9}^{\dagger}$
$\textbf{46.8}^{\dagger}$ $\textbf{46.7}^{\dagger}$
LoTTE-Wri. 41.3 52.7
$\textbf{53.0}^{\dagger}$
$\textbf{53.8}^{\dagger\ddagger}$
$\textbf{53.4}^{\dagger}$
$\textbf{54.1}^{\dagger\ddagger}$
$52.0^{\dagger}$ $\textbf{52.9}^{\dagger}$
LoTTE-Tec. 25.3 36.8
$\textbf{38.2}^{\dagger}$
$\textbf{37.6}^{\dagger}$
$\textbf{37.4}^{\dagger}$
$36.8^{\dagger}$
$36.8^{\dagger}$ $36.8^{\dagger}$
@@ 4.2 Results
Method ($\rightarrow$)
0-shot
k=0
k=1
k=2
k=3
k=4
k=6
k=8
k=10

### v2

@@ Abstract
In Information Retrieval, and more generally in Natural Language Processing, adapting models to specific domains is conducted through fine-tuning. Despite the successes achieved by this method and its versatility, the need for human-curated and labeled data makes it impractical to transfer to new tasks, domains, and/or languages when training data doesn’t exist. Using the model without training (zero-shot) is another option that however suffers an effectiveness cost, especially in the case of first-stage retrievers. Numerous research directions have emerged to tackle these issues, most of them in the context of adapting to a task or a language. However, the literature is scarcer for domain (or topic) adaptation. In this paper, we address this issue of cross-topic discrepancy for a sparse first-stage retriever by transposing a method initially designed for language adaptation. By leveraging pre-training on the target data to learn domain-specific knowledge, this technique alleviates the need for annotated data and expands the scope of domain adaptation. Despite their relatively good generalization ability, we show that even sparse retrievers can benefit from our simple domain adaptation method.
@@ 4.1 Datasets, baselines, and ablations
Dataset ($\downarrow$) BM25 0-shot 0 layer 1 layer 2 layers 4 layers w/o pre-training w/o source
TREC-COVID 58.1 71.3 $68.8^{\dagger}$ $70.5^{\dagger}$ $\textbf{72.1}^{\dagger}$ $70.9^{\dagger}$ $67.9^{\dagger}$ $67.7^{\dagger}$
NFCorpus 24.5 24.9 25.6 25.4 25.1 24.1 25.1 25.6
BioASQ 52.3 43.1 $\textbf{45.4}^{\ddagger}$ 44.2 $\textbf{45.2}^{\ddagger}$ $\textbf{46.2}^{\ddagger}$ $\textbf{44.8}^{\ddagger}$ $\textbf{45.3}^{\ddagger}$
FiQA-2018 23.6 27.3 $27.2^{\dagger}$ $\textbf{27.5}^{\dagger}$ $\textbf{29.5}^{\dagger\ddagger}$ $\textbf{29.0}^{\dagger\ddagger}$ $25.2^{\dagger}$ $\textbf{28.3}^{\dagger}$
TREC-NEWS 31.4 32.1 33.3 32.8 33.0 $\textbf{36.5}^{\ddagger}$ 31.9 33.1
Robust04 40.8 39.6 $\textbf{42.9}^{\ddagger}$ $\textbf{42.9}^{\ddagger}$ $\textbf{43.5}^{\ddagger}$ $\textbf{42.6}^{\ddagger}$ 41.2 $\textbf{42.1}^{\ddagger}$
ANTIQUE 45.4 43.3 42.9 44.0 43.7 44.1 $\textbf{45.0}^{\ddagger}$ $\textbf{45.0}^{\ddagger}$
LoTTE-Rec. 39.3 46.2 $\textbf{47.6}^{\dagger\ddagger}$ $\textbf{47.7}^{\dagger\ddagger}$ $\textbf{47.8}^{\dagger\ddagger}$ $\textbf{46.9}^{\dagger}$ $\textbf{46.8}^{\dagger}$ $\textbf{46.7}^{\dagger}$
LoTTE-Wri. 41.3 52.7 $\textbf{53.0}^{\dagger}$ $\textbf{53.8}^{\dagger\ddagger}$ $\textbf{53.4}^{\dagger}$ $\textbf{54.1}^{\dagger\ddagger}$ $52.0^{\dagger}$ $\textbf{52.9}^{\dagger}$
LoTTE-Tec. 25.3 36.8 $\textbf{38.2}^{\dagger}$ $\textbf{37.6}^{\dagger}$ $\textbf{37.4}^{\dagger}$ $36.8^{\dagger}$ $36.8^{\dagger}$ $36.8^{\dagger}$
@@ 4.2 Results
Method ($\rightarrow$) 0-shot k=0 k=1 k=2 k=3 k=4 k=6 k=8 k=10
@@ 6 Ackowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014444 made by GENCI.